# FedFACT: A Provable Framework for Controllable Group-Fairness Calibration in Federated Learning

### v1

@@ 4.1 Federated Bayes-Optimal Fair Classifier
To investigate the optimal classifier with the group fairness guarantee within FL, we consider the situation that there is a unified fairness constraint at the global level, and each client has additional local fairness restrictions in response to personalized demands. Therefore, it is appropriate to consider a personalized federated model to minimize classification risk and ensure both local and global fairness. Denoting the set of classifiers as =$\mathcal{H}=\{h:\mathcal{X}\to\Delta_{m}\}$, the federated Bayes-optimal fair classification problem can be formulated as
$\min_{\mathbf{h}\in\mathcal{H}^{N}}\mathcal{R}(\mathbf{h}),\quad\text{ s.t. }|\mathscr{D}^{g}(\mathbf{h})|\leq\xi^{g},\quad|\mathscr{D}^{l,k}(\mathbf{h})|\leq\xi^{l,k},\quad k\in[N],$ (1)
where $\xi^{l,k},\xi^{g}$ are positive bounds, $\mathscr{D}^{g}(\mathbf{h}):=\{\mathscr{D}^{g}_{u_{g}}(\mathbf{h})\}_{u_{g}\in\mathcal{U}_{g}}$, $\mathscr{D}^{l,k}(\mathbf{h}):=\{\mathscr{D}^{l,k}_{u_{l,k}}(\mathbf{h})\}_{u_{l,k}\in\mathcal{U}_{l,k}}$, and the inequality applies element-wise. FL model $\mathbf{h}=(h_{1},\dots,h_{N})$ comprises $N$ local classifiers.
@@
If (1) is feasible for any positive $\xi^{l,k}$ and $\xi^{g}$, the client-wise classifier $h^{*}_{k}$ in federated Bayes-optimal fair classifier $\mathbf{h}^{*}=(h^{*}_{1},\dots,h^{*}_{N})$ can be expressed as the linear combination of some deterministic classifiers $\{h^{\prime}_{k,i}\}_{i=1}^{d_{k}}$, i.e., $h_{k}^{*}(x)=\sum_{i=1}^{d_{k}}\alpha_{k,i}h^{\prime}_{k,i}(x),\ \alpha_{k}\in\Delta_{d_{k}}$.
@@ 4.2 In-processing Fair Federated Training via Cost Sensitive Learning
$\mathcal{R}(\mathbf{h})+(\lambda^{(1)}-\lambda^{(2)})^{\top}\mathscr{D}^{g}(\mathbf{h})-(\lambda^{(1)}+\lambda^{(2)})^{\top}\xi^{g}+\sum_{k=1}^{N}(\mu^{(1)}_{k}-\mu^{(2)}_{k})^{\top}\mathscr{D}^{l,k}(\mathbf{h})-(\mu^{(1)}_{k}+\mu^{(2)}_{k})^{\top}\xi^{l,k},$
where $\lambda=\{\lambda^{(1)},\lambda^{(2)}\}$ and $\mu=\{\mu^{(1)}_{k},\mu^{(2)}_{k}\}_{k=1}^{N}$ are the dual parameters. Let $\Lambda:=\{\lambda\in\mathbb{R}^{2|\mathcal{U}_{g}|}_{\geq 0}:\|\lambda\|_{1}\leq B_{d}\}$ and :=$\mathcal{M}:=\{\mu\in\mathbb{R}^{2\sum_{k}|\mathcal{U}_{l,k}|}_{\geq 0}:\|\mu\|_{1}\leq B_{d}\}$. Since $\mathbf{h}$ is random classifier, by Sion’s minimax theorem Variational-Analysis-RockWets98 , the primal problem can be written as a saddle-point optimization
@@
where $\mathbf{M}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{p_{a,k}}\Big{[}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}-\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}-\mu^{(2)}_{k,u_{l,k}})\mathbf{D}^{a,k}_{u_{l,k}}\Big{]}$.
@@
In practice, $\mathbf{s}$ is parameterized by $\phi_{k}$ for $k\in[N]$, formulating the personalized optimization objective $L_{k}(f_{\phi_{k}})=\sum_{i=1}^{n_{k}}\ell_{k}(y_{k,i},\mathbf{s}(x_{k,i};\phi_{k}),a_{k,i})]$ in the FL setting. Appendix B.3 provides the proof of Proposition 3 and further presents that the loss $\ell_{k}$ in (4) is also calibrated for the unified federated Bayes-optimal fair classifier. Inspired by this property, we propose an efficient in-processing algorithm for group-fair classification within FL, as detailed in Algorithm 1. At each iteration $t$, the personalized classifier $h^{t}_{k}$ is obtained by ensembling the unified model $\theta^{t}$ with the local model $\phi^{t}_{k}$. The ensemble weight $w^{t}_{k}$ and its update rule balance the contributions of the unified and local models. The following theorem establishes the personalized regret bound w.r.t. the best model parameter, and further demonstrates that our algorithm achieves an $\epsilon$-approximate stochastic saddle point.
Input : Datasets $\mathcal{D}_{k}=\{x_{k,i},y_{k,i},a_{k,i}\}_{i=1}^{N_{k}}$ from client $k$, $k\in\mathcal{K}$; Communication round $T$; Local round $R$; Initial parameters ${\lambda}^{0},\mu^{0},\theta^{0},\phi^{0}_{k},w^{0}_{k},k\in\mathcal{K}$; Corresponding learning rate $\{\eta,\eta_{d},\eta_{w}\}_{t=1}^{T}$;
@@
         $\mu^{(i),t+1}_{k,u_{l,k}}=\Pi_{\mathcal{M}}\big{[}(3-2i)\sum_{a\in\mathcal{A}}\big{\langle}\widehat{\mathbf{D}}^{a,k}_{u_{l,k}},\widehat{\mathbf{C}}^{a,k}(h_{k}^{t})\big{\rangle}-\xi^{l,k}\big{]},\ i\in[1,2],u_{l,k}\in\mathcal{U}_{l,k}$;
@@
Under mild assumptions, there exists constants $B_{k},B_{L}$, such that the following cumulative regret upper bound is guaranteed for the ensemble personalized models:
@@
Furthermore, suppose that personalized models achieve a $\rho_{t}$-approximate optimal response at iteration $t$, namely ^$\widehat{\mathcal{L}}(\mathbf{h}^{t},\lambda^{t},\mu^{t})\leq\min_{\mathbf{h}}\widehat{\mathcal{L}}(\mathbf{h},\lambda^{t},\mu^{t})+\rho_{t}$, denoting $\bar{\rho}=\sum_{t=1}^{T}\rho_{t}/T$, then the sequences of model and bounded dual parameters comprise an approximate mixed Nash equilibrium:
The complete Theorem 4 with its proof is provided in the Appendix B.4. The regret bound yields a convergence rate of $\mathcal{O}(1/\sqrt{T})$ by appropriately choosing the learning rate, reflecting the stability of the proposed algorithm. Moreover, as $\rho_{t}$ decreases with $t$, the algorithm will gradually converge to the optimal equilibrium. For instance, if $\rho_{t}\propto C/\sqrt{t}$, $\epsilon$ will also exhibit an $\mathcal{O}(1/\sqrt{T})$ convergence rate. The generalization error between the optimal solutions of the empirical dual and primal problems under finite samples is given in Appendix B.5.
Theorem 4 indicates that the weighted classifier $\overline{\mathbf{h}}$ in the output of Algorithm 1 with average $\bar{\lambda},\bar{\mu}$ constructs the approximate saddle point due to the linearity of empirical Lagrange $\widehat{\mathcal{L}}$ to $(\mathbf{h},\lambda,\mu)$. In practice, the classifier weighted by the last few iterations, even only $\bar{\mathbf{h}}^{T}$, delivers a superior accuracy-fairness trade-off, as shown in our experiments.
@@ 4.3 Label-Free Federated Post-Fairness Calibration based on Plug-In Approach
Under Assumption 1, suppose that the primal problem (1) is feasible for any $\xi^{g},\xi^{l,k}>0$ and all non-zero columns of each $\mathbf{D}^{a,k}_{u}$ are distinct, the attribute-aware personalized classifier $\{h^{*}_{k}\}_{k\in[N]}$ is Bayes-optimal under local and global fairness constraints, if $h^{*}_{k}(x,a)=e_{y},y\in\arg\max_{j\in[m]}\Big{(}\big{[}\mathbf{M}^{\lambda^{*},\mu^{*}}(a,k)\big{]}^{\top}\eta(x,a,k)\Big{)}_{j}$, where the dual parameters are determined from $(\lambda^{*},\mu^{*})\in\arg\min_{\lambda,\mu\geq 0}{H(\lambda,\mu)}$,
$H(\lambda,\mu)=$ $\sum_{k\in[N]}\sum_{a\in\mathcal{A}}p_{a,k}\underset{X\sim\mathcal{P}^{X}_{a,c}}{\mathbb{E}}\Big{[}\max_{y\in[m]}\Big{(}\big{[}\mathbf{M}^{\lambda,\mu}(a,k)\big{]}^{\top}\eta(X,a,k)\Big{)}_{y}\Big{]}+\xi^{g}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})$
$\qquad+\sum_{k\in[N]}\xi^{l,k}\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}+\mu^{(2)}_{k,u_{l,k}}).$ (7)
@@
where $\widehat{H}_{k}(\lambda,\mu_{k}):=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\max_{y\in[m]}\big{(}\big{[}\mathbf{M}^{\lambda,\mu}(a_{i},k)\big{]}^{\top}\eta(x_{i},a_{i},k)\big{)}_{y}+\xi^{g}\|\lambda\|_{1}+\frac{\xi^{l,k}}{\hat{p}_{k}}\|\mu_{k}\|_{1}$ is the plug-in estimation of (5). Considering that the non-smoothness of the optimization objective may lead to convergence issues in federated optimization fedcompo-yuan2021federated , we replace the maximum operation in $\widehat{H}_{k}(\lambda,\mu_{k})$ with soft-max weight function $\sigma_{\beta}(x)=\sum_{i=1}^{m}\frac{\exp(x_{i}/\beta)}{\sum_{j=1}^{m}\exp(x_{j}/\beta)}x_{i}$, which reduces to the hard-maximum if temperature $\beta\to 0$. Denoting the relaxed local objective as $\widehat{H}^{\prime}_{k}(\lambda,\mu_{k})$, we propose Algorithm 2 to solve the federated Bayes-optimal fair classifier.
@@
Return Classifiers $\{h_{1},\dots,h_{N}\}$, $h_{k}(x,a):=\arg\max_{j\in[m]}\big{(}\big{[}\widehat{\mathbf{M}}^{\lambda^{T},\mu^{T}}(a,k)\big{]}^{\top}\widehat{\eta}(x,a,k)\big{)}_{j}$;
@@ 4.4 Discussion
In- versus post-processing. In- and post-processing interventions play complementary roles: the former removes bias in representations during training (incurring higher computational cost), and the latter adjusts fairness on model outputs with low overhead, unable to debias learned representations. Their joint deployment supports adaptable fairness calibration in resource-limited, heterogeneous FL.
@@ 5 Experiments
Partition Mehod Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$ Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$ Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$ Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$
$\gamma=0.5$ FedAvg 69.73 0.2766 0.3590 84.52 0.1765 0.2310 89.14 0.1435 0.1308 72.51 0.2158 0.2190
FairFed 59.39 0.1008 0.1022 80.73 0.0983 0.1434 81.85 0.0704 0.1058 64.93 0.0978 0.1728
FedFB 58.09 0.0879 0.0983 81.85 0.0751 0.1165 85.32 0.1188 0.0949 68.91 0.0815 0.1645
FCFL 56.53 0.0646 0.0614 81.91 0.0845 0.1455 83.83 0.0704 0.1090 66.19 0.0749 0.0891
praFFed 59.93 0.0824 0.0968 80.96 0.0591 0.0763 85.45 0.0731 0.0862 66.41 0.0750 0.0818
Cost 64.51 0.0585 0.0941 81.09 0.0262 0.0590 85.60 0.0314 0.0577 69.04 0.0473 0.0689
[…]
Hetero FedAvg 69.12 0.2513 0.4044 85.34 0.1596 0.2358 89.85 0.1360 0.1742 72.36 0.1816 0.3071
FairFed 61.87 0.1825 0.2448 81.85 0.1074 0.1283 82.50 0.0672 0.1415 63.80 0.0924 0.2411
FedFB 60.16 0.1284 0.1332 81.14 0.0949 0.1005 86.00 0.1121 0.1284 68.77 0.0671 0.2316
FCFL 59.96 0.1498 0.1507 79.10 0.0528 0.0596 84.50 0.0657 0.1458 66.05 0.1041 0.1253
praFFed 60.42 0.0902 0.1062 80.12 0.0523 0.0606 85.13 0.0592 0.1152 65.97 0.0946 0.1155
Cost 63.01 0.0773 0.1044 81.04 0.0286 0.0567 86.28 0.0495 0.0761 66.90 0.0202 0.0724
[…]

### v2

@@ 4.1 Federated Bayes-Optimal Fair Classifier
To investigate the optimal classifier with the group fairness guarantee within FL, we consider the situation that there is a unified fairness constraint at the global level, and each client has additional local fairness restrictions in response to personalized demands. Therefore, it is appropriate to consider a personalized federated model to minimize classification risk and ensure both local and global fairness. Denoting the set of classifiers as $\mathcal{H}=\{h:\mathcal{X}\to\Delta_{m}\}$, the federated Bayes-optimal fair classification problem can be formulated as
$\min_{\mathbf{h}\in\mathcal{H}^{N}}\mathcal{R}(\mathbf{h}),\quad\text{ s.t. }|\mathscr{D}^{g}(\mathbf{h})|\leq\xi^{g},\quad|\mathscr{D}^{k}(\mathbf{h})|\leq\xi^{k},\quad k\in[N],$ (1)
where $\xi^{k},\xi^{g}$ are positive bounds, $\mathscr{D}^{g}(\mathbf{h}):=\{\mathscr{D}^{g}_{u_{g}}(\mathbf{h})\}_{u_{g}\in\mathcal{U}_{g}}$, $\mathscr{D}^{k}(\mathbf{h}):=\{\mathscr{D}^{k}_{u_{k}}(\mathbf{h})\}_{u_{k}\in\mathcal{U}_{k}}$, and the inequality applies element-wise. FL model $\mathbf{h}=(h_{1},\dots,h_{N})$ comprises $N$ local classifiers.
@@
If (1) is feasible for any positive $\xi^{k}$ and $\xi^{g}$, the client-wise classifier $h^{*}_{k}$ in federated Bayes-optimal fair classifier $\mathbf{h}^{*}=(h^{*}_{1},\dots,h^{*}_{N})$ can be expressed as the linear combination of some deterministic classifiers $\{h^{\prime}_{k,i}\}_{i=1}^{d_{k}}$, i.e., $h_{k}^{*}(x)=\sum_{i=1}^{d_{k}}\alpha_{k,i}h^{\prime}_{k,i}(x),\ \alpha_{k}\in\Delta_{d_{k}}$.
@@ 4.2 In-processing Fair Federated Training via Cost Sensitive Learning
$\mathcal{R}(\mathbf{h})+(\lambda^{(1)}-\lambda^{(2)})^{\top}\mathscr{D}^{g}(\mathbf{h})-(\lambda^{(1)}+\lambda^{(2)})^{\top}\xi^{g}+\sum_{k=1}^{N}(\mu^{(1)}_{k}-\mu^{(2)}_{k})^{\top}\mathscr{D}^{k}(\mathbf{h})-(\mu^{(1)}_{k}+\mu^{(2)}_{k})^{\top}\xi^{k},$
where $\lambda=\{\lambda^{(1)},\lambda^{(2)}\}$ and $\mu=\{\mu^{(1)}_{k},\mu^{(2)}_{k}\}_{k=1}^{N}$ are the dual parameters. Let $\Lambda:=\{\lambda\in\mathbb{R}^{2|\mathcal{U}_{g}|}_{\geq 0}:\|\lambda\|_{1}\leq B_{d}\}$ and $\mathcal{M}:=\{\mu\in\mathbb{R}^{2\sum_{k}|\mathcal{U}_{k}|}_{\geq 0}:\|\mu\|_{1}\leq B_{d}\}$. Since $\mathbf{h}$ is random classifier, by Sion’s minimax theorem Variational-Analysis-RockWets98, the primal problem can be written as a saddle-point optimization
@@
where $\mathbf{M}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{p_{a,k}}\Big[\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}-\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}-\mu^{(2)}_{k,u_{k}})\mathbf{D}^{a,k}_{u_{k}}\Big]$.
@@
In practice, $\mathbf{s}$ is parameterized by $\phi_{k}$ for $k\in[N]$, formulating the personalized optimization objective $L_{k}(f_{\phi_{k}})=\sum_{i=1}^{n_{k}}\ell_{k}(y_{k,i},\mathbf{s}(x_{k,i};\phi_{k}),a_{k,i})$ in the FL setting. Appendix B.3 provides the proof of Proposition 3 and further presents that the loss $\ell_{k}$ in (4) is also calibrated for the unified federated Bayes-optimal fair classifier. Inspired by this property, we propose an efficient in-processing algorithm for group-fair classification within FL, as detailed in Algorithm 1.
Input : Datasets $\{x_{k,i},y_{k,i},a_{k,i}\}_{i=1}^{N_{k}}$ from client $k$, $k\in\mathcal{K}$; Communication round $T$; Local round $R$; Initial parameters ${\lambda}^{0},\mu^{0},\theta^{0},\phi^{0}_{k},w^{0}_{k},k\in\mathcal{K}$; Learning rate $\{\eta,\eta_{d},\eta_{w}\}_{t=1}^{T}$;
@@
     $\mu^{(i),t+1}_{k,u_{k}}=\Pi_{\mathcal{M}}\big[(3-2i)\sum_{a\in\mathcal{A}}\big\langle\widehat{\mathbf{D}}^{a,k}_{u_{k}},\widehat{\mathbf{C}}^{a,k}(h_{k}^{t})\big\rangle-\xi^{k}\big],\ i\in[1,2],u_{k}\in\mathcal{U}_{k}$;
@@
At each iteration $t$, the personalized classifier $h^{t}_{k}$ is obtained by ensembling the unified model $\theta^{t}$ with the local model $\phi^{t}_{k}$. The ensemble weight $w^{t}_{k}$ and its update rule balance the contributions of the unified and local models. The following theorem establishes the personalized regret bound w.r.t. the best model parameter, and further demonstrates that our algorithm achieves an $\epsilon$-approximate stochastic saddle point.
@@
Under mild assumptions, there exist constants $B_{k},B_{L}$, such that the following cumulative regret upper bound is guaranteed for the ensemble personalized models:
@@
Furthermore, suppose that personalized models achieve a $\rho^{t}$-approximate optimal response at iteration $t$, namely $\widehat{\mathcal{L}}(\mathbf{h}^{t},\lambda^{t},\mu^{t})\leq\min_{\mathbf{h}}\widehat{\mathcal{L}}(\mathbf{h},\lambda^{t},\mu^{t})+\rho^{t}$, denoting $\bar{\rho}^{T}=\sum_{t=1}^{T}\rho^{t}/T$, then the sequences of model and bounded dual parameters comprise an approximate mixed Nash equilibrium:
$\max_{\lambda^{*},\mu^{*}}\frac{1}{T}\sum_{t=1}^{T}\widehat{\mathcal{L}}\left(\mathbf{h}^{t},\lambda^{*},\mu^{*}\right)-\inf_{\mathbf{h}^{*}}\frac{1}{T}\sum_{t=1}^{T}\widehat{\mathcal{L}}\left(\mathbf{h}^{*},\lambda^{t},\mu^{t}\right)\leq\epsilon^{T}:=\bar{\rho}^{T}+16B_{d}^{2}\sqrt{\frac{1}{T}}.$ (6)
The complete Theorem 4 with its proof is provided in the Appendix B.4. The regret bound yields a convergence rate of $\mathcal{O}(1/\sqrt{T})$ by appropriately choosing the learning rate, reflecting the stability of the proposed algorithm. Moreover, as $\rho^{t}$ decreases with $t$, the algorithm will gradually converge to the optimal equilibrium. For instance, if $\rho^{t}\propto C/\sqrt{t}$, $\epsilon$ will also exhibit an $\mathcal{O}(1/\sqrt{T})$ convergence rate. The generalization error between the optimal solutions of the empirical dual and primal problems under finite samples is given in Appendix B.5.
@@ 4.3 Label-Free Federated Post-Fairness Calibration based on Plug-In Approach
Under Assumption 1, suppose that the primal problem (1) is feasible for any $\xi^{g},\xi^{k}>0$ and all non-zero columns of each $\mathbf{D}^{a,k}_{u}$ are distinct, the attribute-aware personalized classifier $\{h^{*}_{k}\}_{k\in[N]}$ is Bayes-optimal under local and global fairness constraints, if $h^{*}_{k}(x,a)=e_{y},y\in\arg\max_{j\in[m]}\Big(\big[\mathbf{M}^{\lambda^{*},\mu^{*}}(a,k)\big]^{\top}\eta(x,a,k)\Big)_{j}$, where the dual parameters are determined from $(\lambda^{*},\mu^{*})\in\arg\min_{\lambda,\mu\geq 0}{H(\lambda,\mu)}$,
$H(\lambda,\mu)=$ $\sum_{k\in[N]}\sum_{a\in\mathcal{A}}p_{a,k}\underset{X\sim\mathcal{P}^{X}_{a,c}}{\mathbb{E}}\Big[\max_{y\in[m]}\Big(\big[\mathbf{M}^{\lambda,\mu}(a,k)\big]^{\top}\eta(X,a,k)\Big)_{y}\Big]+\xi^{g}\|\lambda\|_{1}+\sum_{k\in[N]}\xi^{k}\|\mu_{k}\|_{1}.$ (7)
@@
where $\widehat{H}_{k}(\lambda,\mu_{k}):=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\max_{y\in[m]}\big(\big[\mathbf{M}^{\lambda,\mu}(a_{i},k)\big]^{\top}\eta(x_{i},a_{i},k)\big)_{y}+\xi^{g}\|\lambda\|_{1}+\frac{\xi^{k}}{\hat{p}_{k}}\|\mu_{k}\|_{1}$ is the plug-in estimation of (7). Considering that the non-smoothness of the optimization objective may lead to convergence issues in federated optimization fedcompo-yuan2021federated, we replace the maximum operation in $\widehat{H}_{k}(\lambda,\mu_{k})$ with soft-max weight function $\sigma_{\beta}(x)=\sum_{i=1}^{m}\frac{\exp(x_{i}/\beta)}{\sum_{j=1}^{m}\exp(x_{j}/\beta)}x_{i}$, which reduces to the hard-maximum if temperature $\beta\to 0$. Denoting the relaxed local objective as $\widehat{H}^{\prime}_{k}(\lambda,\mu_{k})$, we propose Algorithm 2 to solve the federated Bayes-optimal fair classifier.
@@
Return Classifiers $\{h_{1},\dots,h_{N}\}$, $h_{k}(x,a):=\arg\max_{j\in[m]}\big(\big[\widehat{\mathbf{M}}^{\lambda^{*},\mu^{*}}(a,k)\big]^{\top}\widehat{\eta}(x,a,k)\big)_{j}$;
@@ 4.4 Discussion
In- versus post-processing. In- and post-processing interventions play complementary roles: the former removes bias in representations during training (incurring higher computational cost), and the latter adjusts fairness on model outputs with low overhead, unable to debias learned representations. Both of them support adaptable fairness calibration in resource-limited, heterogeneous FL. Note that combining the in-processing and post-processing methods is theoretically unjustifiable, as the in-processing classifier is not designed to approximate the Bayes score function.
@@ 5 Experiments
Partition Method Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$ Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$ Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$ Acc $\mathscr{D}^{global}$ $\mathscr{D}^{local}$
$\gamma=0.5$ FedAvg 69.73 0.2766 0.3590 84.52 0.1765 0.2310 89.14 0.1435 0.1308 67.56 0.2620 0.2462
FairFed 59.39 0.1008 0.1022 80.73 0.0983 0.1434 81.85 0.0704 0.1058 60.99 0.1165 0.1733
FedFB 58.09 0.0879 0.0983 81.85 0.0751 0.1165 85.32 0.1188 0.0949 64.35 0.0814 0.1326
FCFL 56.53 0.0646 0.0614 81.91 0.0845 0.1455 83.83 0.0704 0.1090 61.07 0.1260 0.1189
praFFed 59.93 0.0824 0.0968 80.96 0.0591 0.0763 85.45 0.0731 0.0862 62.60 0.0736 0.0806
Cost 64.51 0.0585 0.0941 81.09 0.0262 0.0590 85.60 0.0314 0.0577 65.79 0.0487 0.0674
FedFACTgg (In) 61.19 0.0344 0.0761 82.05 0.0015 0.0408 86.49 0.0205 0.0544 63.79 0.0493 0.0568
FedFACTll (In) 61.81 0.0600 0.0636 82.44 0.0140 0.0508 85.90 0.0461 0.0312 63.93 0.0434 0.0487
FedFACTg&lg\&l (In) 61.17 0.0407 0.0732 82.04 0.0014* 0.0401 86.15 0.0382 0.0473 62.51 0.0366 0.0413
FedFACTgg (Post) 67.27 0.0128* 0.0660 82.83* 0.0173 0.0276 87.25 0.0089* 0.0253 66.15 0.0175 0.0181*
FedFACTll (Post) 67.49* 0.0315 0.0552* 82.79 0.0154 0.0267* 87.36* 0.0127 0.0163* 66.54* 0.0197 0.0240
FedFACTg&lg\&l (Post) 67.33 0.0139 0.0641 82.74 0.0134 0.0274 87.06 0.0093 0.0172 66.52 0.0162* 0.0184
Hetero FedAvg 69.12 0.2513 0.4044 85.34 0.1596 0.2358 89.85 0.1360 0.1742 67.45 0.2037 0.3068
FairFed 61.87 0.1825 0.2448 81.85 0.1074 0.1283 82.50 0.0672 0.1415 59.81 0.0949 0.1624
FedFB 60.16 0.1284 0.1332 81.14 0.0949 0.1005 86.00 0.1121 0.1284 63.68 0.0780 0.2039
FCFL 59.96 0.1498 0.1507 79.10 0.0528 0.0596 84.50 0.0657 0.1458 61.05 0.1134 0.1425
praFFed 60.42 0.0902 0.1062 80.12 0.0523 0.0606 85.13 0.0592 0.1152 60.84 0.0932 0.1246
Cost 63.01 0.0773 0.1044 81.04 0.0286 0.0567 86.28 0.0495 0.0761 62.11 0.0315 0.0730
FedFACTgg (In) 60.33 0.0665 0.0841 82.09* 0.0122 0.0962 86.18 0.0188 0.0731 62.05 0.0380 0.0577
FedFACTll (In) 60.22 0.0730 0.0753 81.19 0.0208 0.0250 86.58 0.0424 0.0426 61.96 0.0471 0.0473
FedFACTg&lg\&l (In) 61.44 0.0676 0.0789 81.19 0.0055 0.0239* 86.38 0.0355 0.0634 61.48 0.0322 0.0364
FedFACTgg (Post) 64.36 0.0398* 0.0699 81.09 0.0047* 0.0257 87.95 0.0094 0.0344 65.32 0.0199* 0.0343
FedFACTll (Post) 64.38 0.0479 0.0740 81.27 0.0049 0.0306 88.05* 0.0112 0.0217* 65.68* 0.0246 0.0120*
FedFACTg&lg\&l (Post) 64.41* 0.0408 0.0680* 81.31 0.0053 0.0293 87.75 0.0088* 0.0247 65.19 0.0201 0.0131